# Complete complex Finsler metrics and uniform equivalence of the Kobayashi metric

### v3

@@ 7 Uniform equivalence of Kobayashi metric
Suppose that $M$ is a Lempert manifold. And let $(M,h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $-A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then Kobayashi metric $\mathfrak{K}_{M}$ satisfies
@@
Here $C>0$ is only depending only $A,B$, and independent of complex dimension $n$ of complex manifold $M$.
@@
$h(z,v)=(\phi^{*}h)(0,\xi)\leq\frac{4}{B}\omega_{\mathcal{P}}(0,\xi)=\frac{4}{B}|\xi|^{2}_{\mathbb{C}}.$ (7.1)
Hence, $|\xi|^{2}_{\mathbb{C}}\geq\frac{B}{4}h(z,v)$. By the definition of Kobayashi metric, we obtain $\mathfrak{K}^{2}_{M}(z,v)\geq\frac{B}{4}h(z,v)$.
@@
Combing Theorems 5.1, 5.2 and Theorem 7.1, we prove the following corollary.
@@
Suppose that $\mathcal{D}$ is a bounded strongly convex domain with smooth boundary in $\mathbb{C}^{n}$. And let $(\mathcal{D},h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $-A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then Kobayashi metric and Carathéodory metric satisfies
@@
Here $C>0$ is only depending only $A,B$, and independent of complex dimension $n$ of complex manifold $\mathcal{D}$.
@@
For a suitable choice of positive constant $C$, suppose that $(M,h)$ is a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $-A\leq K_{h}\leq B$ for some positive constants $A$ and $B$. If Carathédory metric $\mathfrak{C}^{2}$ on complex manifold $M$ is a smooth strongly pseudoconvex complex Finsler mtric. Then $H=C\mathfrak{C}^{2}+h$ and the Kähler metric $h$ are equivalent. What’s more, $M$ is a complete Kobayashi-hyperbolic manifold.
@@
By Wong’s result [27], we know that the holomorphic sectional curvature of the Carathédory metric is less than $-4$. In the similar way as Theorem 6.4, we know that the holomorphic sectional curvature of $H$ is bounded above by a negative constant $-\frac{4}{C}+B(0<C<\frac{4}{B})$. By assumptions and Theorem 4.1, applying Theorem 6.1 to the case where a holomorphic map $f:M\rightarrow M$ is an identity map $id$, we have
@@
Acknowledgement: The author thanks the referees for carefully reading the manuscript and their valuable corrections/suggestions which improved the presentation of the paper. This work was supported by National Natural Science Foundation of China (Grant No. 12401101, No. 12461008, No. 12461013) and Natural Science Foundation of Jiangxi Provence in China (Grant No. 20232ACB201005).

### v4

@@ 4 Uniform equivalence of Kobayashi metric
Suppose that $M$ is a Lempert manifold. And let $(M,h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then Kobayashi metric $\mathfrak{K}_{M}$ satisfies
@@
Here $C>0$ is depending only $A,B$, and independent of complex dimension $n$ of complex manifold $M$.
@@
$h(z,v)=(\phi^{*}h)(0,\xi)\leq\frac{4}{B}\omega_{\mathcal{P}}(0,\xi)=\frac{4}{B}|\xi|^{2}_{\mathbb{C}}.$
Hence, $|\xi|^{2}_{\mathbb{C}}\geq\frac{B}{4}h(z,v)$. By the definition of Kobayashi metric, we obtain
$\mathfrak{K}^{2}_{M}(z,v)\geq\frac{B}{4}h(z,v).$ (4.1)
@@
Combining Theorems 2.1, 2.2 and Theorem 4.1, we prove the following corollary.
@@
Suppose that $\mathcal{D}$ is a bounded strongly convex domain with smooth boundary in $\mathbb{C}^{n}$. And let $(\mathcal{D},h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then Kobayashi metric and Carathéodory metric satisfies
@@
Here $C>0$ is only depending $A,B$, and independent of complex dimension $n$ of complex manifold $\mathcal{D}$.
@@
For a suitable choice of positive constant $C$, suppose that $(M,h)$ is a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $A\leq K_{h}\leq B$ for some positive constants $A$ and $B$. If Carathéodory metric $\mathfrak{C}^{2}$ on complex manifold $M$ is a smooth strongly pseudoconvex complex Finsler metric. Then $H=C\mathfrak{C}^{2}+h$ and the Kähler metric $h$ are equivalent. What’s more, $M$ is a complete Kobayashi-hyperbolic manifold.
@@
By Wong’s result [28], we know that the holomorphic sectional curvature of the Carathéodory metric is less than 4-4. In the similar way as Theorem 3.4, we know that the holomorphic sectional curvature of $H$ is bounded above by a negative constant $\frac{4}{C}+B(0<C<\frac{4}{B})$. By assumptions and Theorem 2.5, applying Theorem 3.1 to the case where a holomorphic map $f:M\rightarrow M$ is an identity map $id$, we have
@@
Acknowledgement: The author thanks the referees for carefully reading the manuscript and their valuable corrections/suggestions which improved the presentation of the paper. This work was supported by National Natural Science Foundation of China (Grant No. 12401101, No. 12461008, No. 12461013) and Natural Science Foundation of Jiangxi Province in China (Grant No. 20232ACB201005).